OpenStudy (christinepeav):

what are the two arithmetic means between -13 and 8?

OpenStudy (sshayer):

let a1 and a2 be two A.M's between -13 and 8 -13,a1,a2,8 are in A.P. let d be common difference. first term a=-13 n=4 last term =8 \[l=a+(n-1)d\] 8=-13+(4-1)d 8+13=3d d=21/3=7 a1=-13+7=-6 a2=-6+7=?